# Learning Safe Strategies for Value Maximizing Buyers in Uniform Price Auctions

### v2

@@ 2.  Model
Preliminaries. There are $n$ buyers (bidders) indexed by $i\in[n]$, and $K$ identical units of a single good. Each bidder $i$ has a fixed, private valuation curve, denoted by $\mathbf{v}_{i}\in\mathbb{R}^{K}_{+}$ that has diminishing marginal returns property, i.e., $v_{i,1}\geq v_{i,2}\geq\dots\geq v_{i,K}$ which is standard in literature (Brânzei et al., 2023; Goldner et al., 2020). As we study optimal bidding strategies from the perspective of a single bidder, we drop the index $i$ when the context is clear. The maximum total demand for bidder $i$, denoted by $M\in[K]$, is defined as $\min\{j\in[K-1]:v_{j+1}=0\}$. If such an index does not exist, we set $M$ as $K$. Hence, without loss of generality, we assume $\mathbf{v}\in\mathbb{R}_{+}^{M}$. For each $\mathbf{v}=[v_{1},\dots,v_{M}]$, we define the average cumulative valuation vector as $\mathbf{w}=[w_{1},\dots,w_{M}]$,
$\text{where}\quad w_{j}$ $=\frac{1}{j}\sum_{\ell\leq j}v_{\ell},\forall j\in[M]\,.$ (1)
@@ 2.1.  Auction Format and Bidders’ Behavior
Allocation and Payment Rule. In a uniform price auction, each bidder $i$ submits a sorted bid vector $\mathbf{b}$ using $m$-uniform bidding language (see Definition 1). The auctioneer collects the bids (entries of the bid vector) from all the bidders, sorts them in non-increasing order, and allocates units to the bidders with the top $K$ bids (also termed as ‘winning’ bids). That is, if bidder $i$ has $j$ bids in the top $K$ positions, they are allocated $j$ units. For ease of exposition, we assume there are no ties (or ties are always broken in the favor of the bidder in consideration). See Appendix F for our discussion on handling ties. We assume the auction follows the last-accepted-bid payment rule (bidders pay the $K^{th}$ highest bid per unit) which is widely used for uniform price auctions in practice (Regulations, 2019; Garbade and Ingber, 2005).
Let $\bm{\beta}_{-}$ denote the bids submitted by all bidders except bidder $i$ and $\bm{\beta}_{-}^{-(j)}$ be the $j^{th}$ smallest among the top $K$ competing bids (i.e., bids from all bidders except bidder $i$). Suppose bidder $i$ submits $\mathbf{b}$, such that the bid profile is $\bm{\beta}:=(\mathbf{b};\bm{\beta}_{-})$. Let $x(\bm{\beta})$ and $p(\bm{\beta})$ denote the number of units allocated to bidder $i$ and the clearing price (i.e., the per-unit price which is the $K^{th}$ highest submitted bid), respectively. The total value obtained by the bidder is $V(\bm{\beta})=\sum_{j\leq x(\bm{\beta})}v_{j},$ while the total payment made is $P(\bm{\beta})=p(\bm{\beta})\cdot x(\bm{\beta}).$
Bidding Language. Multi-unit auctions allocate a large number of identical units, requiring efficient ways for the bidders to express preferences. A common approach is standard bidding, where bidders submit a vector of bids, one for each unit (Brânzei et al., 2023; Galgana and Golrezaei, 2024; Babaioff et al., 2023; Birmpas et al., 2019; Potfer et al., 2024). Although expressive, this becomes computationally impractical when the number of units, $K$, is large, as in EU ETS emission permit auctions and Treasury auctions. To address this, we consider a bidding language called $m$-uniform bidding, for any $m\in\mathbb{N}$ [^4] This format generalizes the uniform bidding format (De Keijzer et al., 2013; Birmpas et al., 2019) and aligns with practical languages like those in product-mix auctions (Klemperer, 2009) and piecewise-linear bidding (Nisan, 2015). In $m$-uniform bidding, bidders submit $m$ bid-quantity pairs $(b_{i},q_{i})$, where $b_{i}$ is the bid value per unit and $q_{i}$ is the quantity demanded:
@@
For a fixed $m\in\mathbb{N}$, a $m$-uniform bidding strategy is characterized by $m$ bid-quantity pairs, denoted as $\mathbf{b}:=\langle(b_{1},q_{1}),\dots,(b_{m},q_{m})\rangle$, where $b_{1}>b_{2}>\dots>b_{m}>0$ and $q_{j}>0$, $j\in[m]$. This $m$-uniform bid can be equivalently expressed as a vector (similar to the standard bidding format) in which the first $q_{1}$ bids are $b_{1}$, followed by $q_{2}$ bids of $b_{2}$, and so on.
We define $\mathbf{b}[1:\ell]=\langle(b_{1},q_{1}),\dots,(b_{\ell},q_{\ell})\rangle$, for all $\ell<m$, to represent the first $\ell$ bid-quantity pairs within a $m$-uniform bidding strategy $\mathbf{b}=\langle(b_{1},q_{1}),\dots,(b_{m},q_{m})\rangle$. We further define $Q_{j}=\sum_{\ell=1}^{j}q_{\ell}$ for all $j\in[m]$ as the total quantity demanded in the first $j$ bid-quantity pairs, with $Q_{0}=0$, where we assume, without loss of generality that $Q_{m}\leq M$ [^5] Suppose the bidder bids for, and wins more than $M$ units. There is no additional value being allocated over $M$ units, but the total payment increases (assuming the clearing price is positive), potentially violating the RoI constraint. If $m=M$, the bidding format is equivalent to standard bidding but in practice, bidders often submit only a few bid-quantity pairs. For instance, in the 2023 EU ETS auctions, bidders submitted $\sim 4.35$ bid-quantity pairs per auction on average (EEX, 2023).
@@
Here, $\gamma$ is defined as the target RoI which is private and fixed. Without loss of generality, we assume $\gamma=0$ (or equivalently the valuation curve $\mathbf{v}$ is scaled by $\frac{1}{1+\gamma}$) for the rest of this work. For $\gamma=0$, the RoI constraint implies the value obtained in an auction is at least the payment. We illustrate how the auction operates in Example 1.
@@
Consider an auction with $n=2$ bidders, and $K=5$ identical units. The valuations are: $\mathbf{v}_{1}=[6,4,3,1,1]$ and $\mathbf{v}_{2}=[5,3,1,1,0]$. Both the bidders are value maximizing buyers with target RoI, $\gamma_{1}=\gamma_{2}=0$. Suppose $m=2$ and the bids submitted by the bidders are $\mathbf{b}_{1}=\langle(5,2),(3,3)\rangle$ and $\mathbf{b}_{2}=\langle(4,2),(2,2)\rangle$. The bid profile: $\bm{\beta}=[5,5,4,4,3,3,3,2,2]$ and top $K=5$ winning bids are $[\underline{5},\underline{5},4,4,\underline{3}]$. Bidder 1 is allocated $3$ units as they have $3$ bids (underlined) in the winning bids, and bidder 2 gets the remaining $2$ units. The clearing price $p(\bm{\beta})=3$, $V_{1}(\bm{\beta})=6+4+3=13,V_{2}(\bm{\beta})=5+3=8,P_{1}(\bm{\beta})=3\cdot 3=9$, and $P_{2}(\bm{\beta})=2\cdot 3=6$. The RoI constraint is satisfied for both the bidders as $13>9$ and $8>6$.
@@ 2.2.  Learning to Bid in Repeated Settings
In practice, most multi-unit auctions, such as emission permit auctions and Treasury auctions, are conducted in a repeated setting. Formally, the auction described in the previous section takes place sequentially over $T$ rounds indexed by $t\in[T]$. We assume that the valuation vector $\mathbf{v}$ is fixed over the $T$ rounds which is a standard assumption in the literature (Brânzei et al., 2023; Galgana and Golrezaei, 2024; Potfer et al., 2024). Similarly, the target RoI is also assumed to be fixed over the $T$ rounds.
We now extend the notations from the previous section to the repeated setting. Formally, in this setting, $\bm{\beta}_{-}^{t}$ denotes the bids submitted by all bidders except bidder $i$ in round $t$ and $\bm{\beta}_{-,t}^{-(j)}$ is the $j^{th}$ smallest among the top $K$ competing bids in round $t$. In round $t$, if bidder $i$ submits a bid $\mathbf{b}^{t}$, the bid profile is $\bm{\beta}^{t}:=(\mathbf{b}^{t};\bm{\beta}_{-}^{t})$. Let $x(\bm{\beta}^{t})$ and $p(\bm{\beta}^{t})$ denote the number of units allocated to bidder $i$ and the clearing price, respectively, in round $t$. The total value obtained by the bidder is $V(\bm{\beta}^{t})=\sum_{j\leq x(\bm{\beta}^{t})}v_{j},$ while the total payment made is $P(\bm{\beta}^{t})=p(\bm{\beta}^{t})\cdot x(\bm{\beta}^{t}).$
RoI Constraints in Repeated Setting. In the repeated setting, we require the bidders to satisfy the RoI constraint described in Eq. 2 in every round. A bidding strategy $\mathbf{b}$ is called feasible for a sequence of competing bids $[\bm{\beta}_{-}^{t}]_{t\in[T]}$, if the RoI constraint is satisfied for every round $t\in[T]$.
@@
Our notion of RoI constraints in the repeated setting aligns with the definitions in Wilkens et al. (2016, 2017); Lv et al. (2023). Similar constraints are considered by Lucier et al. (2024) and Gaitonde et al. (2023), termed as marginal RoI (or value) and maximum bid constraints (up to scaling factors), respectively. Unlike the aggregate constraints over $T$ rounds typically assumed in online ad auction literature (Deng et al., 2021, 2024; Feng et al., 2023; Deng et al., 2023a), we enforce RoI constraints for each auction individually. This distinction reflects the fundamental differences between the two settings: ad auctions often occur simultaneously and frequently, with values in the order of cents, whereas Treasury and emission permit auctions occur sequentially over longer horizons, with units valued in millions. Hence, bidders in such auctions prioritize profitability in each auction rather than waiting for an indefinite period, leading to per-round RoI constraints. [^6] Although EU ETS emission permit auctions are scheduled to occur regularly, regulations stipulate that an auction may be canceled if the bidders’ demand falls short of the supply of permits or if the clearing price of the auction does not meet the reserve price (Regulations, 2019). Hence, the bidders are more likely to ensure RoI feasibility in each round.
@@ 2.3.  Objective and Performance Metric
We consider an online setting where a bidder seeks to maximize cumulative value while satisfying RoI constraints and managing uncertainty about competing bids. In this setting, bidders privately submit their bids, and the auctioneer allocates items and sets prices based on these bids. In round $t$, the learning algorithm maps the feedback information set from the first $t-1$ rounds, denoted by $\mathcal{I}^{t-1}$, to a bid $\mathbf{b}^{t}$, where this mapping may be deterministic or random. Under full information feedback, the information set $\mathcal{I}^{t-1}=(\bm{\beta}^{1},\bm{\beta}^{2},\ldots,\bm{\beta}^{t-1}),$ which includes all bids from previous rounds. In the bandit setting, the information set $\mathcal{I}^{t-1}=(V(\bm{\beta}^{1}),P(\bm{\beta}^{1}),\ldots,V(\bm{\beta}^{t-1}),P(\bm{\beta}^{t-1}))$ includes only the value and price outcomes of the previous rounds.
@@
$\mathsf{REG}=\max_{\mathbf{b}\in\mathscr{B}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})-\sum_{t=1}^{T}\mathbb{E}[V(\mathbf{b}^{t};\bm{\beta}_{-}^{t})],$ (3)
where the expectation is with respect to any randomness in the learning algorithm. Here, $\mathscr{B}$ is the class of bidding strategies (formally characterized in Section 3). We require the bidding strategies in $\mathscr{B}$ to be RoI feasible for the sequence of competing bids $[\bm{\beta}_{-}^{t}]_{t\in[T]}$. In Section 4, we consider the case when the both the clairvoyant and the learner choose strategies from the same class and aim to obtain sublinear regret. Later, we consider more challenging settings where the clairvoyant can choose the optimal strategy from much richer classes of strategies compared to the learner. In these cases, we obtain sublinear $\alpha$-approximate regret, where $\alpha\in(0,1]$ is the richness ratio (see details in Section 5).

### v4

@@ 2. Model
Preliminaries. There are $n$ buyers (bidders) indexed by $i\in[n]$, and $K$ identical units of a single good. Each bidder $i$ has a fixed, private valuation curve, denoted by $\mathbf{v}_{i}\in\mathbb{R}^{K}_{+}$ that has diminishing marginal returns property, i.e., $v_{i,1}\geq v_{i,2}\geq\dots\geq v_{i,K}$ which is standard in literature (Brânzei et al., 2023; Goldner et al., 2020). As we study optimal bidding strategies from the perspective of a single bidder, we drop the index $i$ when the context is clear. The maximum total demand for bidder $i$, denoted by $M\in[K]$, is defined as $\min\{j\in[K-1]:v_{j+1}=0\}$. If such an index does not exist, we set $M$ as $K$. Hence, without loss of generality, we assume $\mathbf{v}\in\mathbb{R}_{+}^{M}$. For each $\mathbf{v}=[v_{1},\dots,v_{M}]$, we define the average cumulative valuation vector as $\mathbf{w}=[w_{1},\dots,w_{M}]$, where
$w_{j}$ $=\frac{1}{j}\sum_{\ell\leq j}v_{\ell},\forall j\in[M]\,.$ (1)
@@ 2.1. Auction Format and Bidders’ Behavior
Allocation and Payment Rule. In a uniform price auction, each bidder $i$ submits a sorted bid vector $\mathbf{b}$ using $m$-uniform bidding language (see Definition 1). The auctioneer collects the bids (entries of the bid vector) from all the bidders, sorts them in non-increasing order, and allocates units to the bidders with the top $K$ bids (also termed as ‘winning’ bids). That is, if bidder $i$ has $j$ bids in the top $K$ positions, they are allocated $j$ units. For ease of exposition, we assume there are no ties (or ties are always broken in the favor of the bidder in consideration). See Appendix G for our discussion on handling ties. We assume the auction follows the last-accepted-bid payment rule (bidders pay the $K^{th}$ highest bid per unit), which is widely used for uniform price auctions in practice (Regulations, 2019; Garbade and Ingber, 2005; Burkett and Woodward, 2020).
Let $\bm{\beta}_{-}$ denote the vector of bids submitted by all bidders except bidder $i$. Let $\textsc{top}_{K}(\bm{\beta}_{-})\subseteq\bm{\beta}_{-}$ be the multiset of the top $K$ highest bids among these competing bids. Furthermore, we define $\bm{\beta}_{-}^{-(j)}$ to be the $j^{\text{th}}$ smallest element in $\textsc{top}_{K}(\bm{\beta}_{-})$. Suppose bidder $i$ submits $\mathbf{b}$, such that the bid profile is $\bm{\beta}:=(\mathbf{b};\bm{\beta}_{-})$. Let $x(\bm{\beta})$ and $p(\bm{\beta})$ denote the number of units allocated to bidder $i$ and the clearing price (i.e., the per-unit price which is the $K^{th}$ highest submitted bid), respectively. The total value obtained by the bidder is $V(\bm{\beta})=\sum_{j\leq x(\bm{\beta})}v_{j},$ while the total payment made is $P(\bm{\beta})=p(\bm{\beta})\cdot x(\bm{\beta}).$
Bidding Language. Multi-unit auctions allocate a large number of identical units, requiring efficient ways for the bidders to express preferences. A common approach is standard bidding, where bidders submit a vector of bids, one for each unit (Brânzei et al., 2023; Galgana and Golrezaei, 2024; Babaioff et al., 2023; Birmpas et al., 2019; Potfer et al., 2024). Although expressive, this becomes computationally impractical when the number of units, $K$, is large, as in EU ETS emission permit auctions and Treasury auctions. To address this, we consider a bidding language called $m$-uniform bidding, for any $m\in\mathbb{N}$. [^9] This format generalizes the uniform bidding format (De Keijzer et al., 2013; Birmpas et al., 2019) and aligns with practical languages like those in product-mix auctions (Klemperer, 2009) and piecewise-linear bidding (Nisan, 2015). In $m$-uniform bidding format, bidders submit $m$ bid-quantity pairs $(b_{i},q_{i})$, where $b_{i}$ is the bid value per unit and $q_{i}$ is the quantity demanded:
@@
For a fixed $m\in\mathbb{N}$, a $m$-uniform bidding strategy is characterized by $m$ bid-quantity pairs, denoted as
$\mathbf{b}:=\langle(b_{1},q_{1}),\dots,(b_{m},q_{m})\rangle\,$,
where $b_{1}>b_{2}>\dots>b_{m}>0$ and $q_{j}>0$, $j\in[m]$. This $m$-uniform bidding strategy can be equivalently expressed as a vector (similar to the standard bidding format) in which the first $q_{1}$ bids are $b_{1}$, followed by $q_{2}$ bids of $b_{2}$, and so on.
We define $\mathbf{b}[1:\ell]=\langle(b_{1},q_{1}),\dots,(b_{\ell},q_{\ell})\rangle$, for all $\ell<m$, to represent the first $\ell$ bid-quantity pairs within a $m$-uniform bidding strategy $\mathbf{b}=\langle(b_{1},q_{1}),\dots,(b_{m},q_{m})\rangle$. We further define
$Q_{j}=\sum_{\ell=1}^{j}q_{\ell},\qquad\text{ for all $j\in[m]$}$
as the total quantity demanded in the first $j$ bid-quantity pairs, with $Q_{0}=0$, and we assume, without loss of generality that $Q_{m}\leq M$. [^10] Suppose the bidder bids for, and wins more than $M$ units. There is no additional value being allocated over $M$ units, but the total payment increases (assuming the clearing price is positive), potentially violating the RoI constraint. If $m=M$, the bidding format is equivalent to standard bidding, but in practice, bidders often submit only a few bid-quantity pairs. For instance, in the 2023 EU ETS auctions, bidders submitted $\sim 4.35$ pairs per auction on average (EEX, 2023).
@@
Here, $\gamma$ is defined as the target RoI which is private and fixed. Without loss of generality, we assume $\gamma=0$ (or equivalently, the valuation curve $\mathbf{v}$ is scaled by $\frac{1}{1+\gamma}$) for the rest of this work. For $\gamma=0$, the RoI constraint implies the value obtained in an auction is at least the payment.
@@
Consider an auction with $n=2$ bidders, and $K=5$ identical units. The valuations are: $\mathbf{v}_{1}=[6,4,3,1,1]$ and $\mathbf{v}_{2}=[5,3,1,1,0]$. Both the bidders are value maximizing buyers with target RoI, $\gamma_{1}=\gamma_{2}=0$. Suppose $m=2$ and the bids submitted by the bidders are $\mathbf{b}_{1}=\langle(5,2),(3,3)\rangle$ and $\mathbf{b}_{2}=\langle(4,2),(2,2)\rangle$. The bid profile: $\bm{\beta}=[5,5,4,4,3,3,3,2,2]$ and top $K=5$ winning bids are $\underline{5},\underline{5},4,4,\underline{3}]$. Bidder 1 is allocated $3$ units as they have $3$ bids (underlined) in the winning bids, and bidder 2 gets the remaining $2$ units. The clearing price $p(\bm{\beta})=3$, $V_{1}(\bm{\beta})=6+4+3=13,V_{2}(\bm{\beta})=5+3=8,P_{1}(\bm{\beta})=3\cdot 3=9$, and $P_{2}(\bm{\beta})=2\cdot 3=6$. The RoI constraint is satisfied for both the bidders as $13>9$ and $8>6$.
@@ 2.2. Learning to Bid in Repeated Settings
In practice, most multi-unit auctions, such as emission permit auctions and Treasury auctions, are conducted in a repeated setting. Formally, the auction described in the previous section takes place sequentially over $T$ rounds indexed by $t\in[T]$. Throughout this work, we assume that $T$ is known to the bidder. [^11] In case $T$ is unknown, the bidder can use the standard doubling trick (Auer et al., 2002; Cesa-Bianchi and Lugosi, 2006) while designing learning algorithms. We assume that the valuation vector $\mathbf{v}$ is fixed over the $T$ rounds which is a standard assumption in the literature (Brânzei et al., 2023; Galgana and Golrezaei, 2024; Potfer et al., 2024). Nonetheless, we discuss the case of time-varying valuation vectors in Section 6.3 and show that the techniques developed for the scenario when the valuation vector is fixed can be leveraged to learn to bid when the valuation vectors change across rounds.
We now extend the notations from the previous section to the repeated setting. Formally, in this setting, $\bm{\beta}_{-}^{t}$ denotes the bids submitted by all bidders except bidder $i$ in round $t$ and $\bm{\beta}_{-,t}^{-(j)}$ is the $j^{th}$ smallest among the top $K$ competing bids in round $t$. In round $t$, if bidder $i$ submits a bid $\mathbf{b}^{t}$, the bid profile is $\bm{\beta}^{t}:=(\mathbf{b}^{t};\bm{\beta}_{-}^{t})$. Let $x(\bm{\beta}^{t})$ and $p(\bm{\beta}^{t})$ denote the number of units allocated to bidder $i$ and the clearing price, respectively, in round $t$. For the $t^{th}$ auction, the total value obtained by the bidder is $V(\bm{\beta}^{t})=\sum_{j\leq x(\bm{\beta}^{t})}v_{j},$ while the total payment made is $P(\bm{\beta}^{t})=p(\bm{\beta}^{t})\cdot x(\bm{\beta}^{t}).$
Nature of Competing Bids. We model the competing bids in each round as being chosen adversarially to capture realistic scenarios where the bidder has partial or no information about their competitors’ strategy. We assume that the competing bids are selected by an oblivious adversary, i.e., the bids in round $t$ are chosen adversarially but independently of the bidder’s strategies in earlier rounds $\tau=1,\dots,t-1$. This is equivalent to the case where all competing bids are fixed in advance, prior to the first round, and without any knowledge of the bidder’s realized strategies. In our setting, an oblivious adversary models a sufficiently large market where the bids submitted by a single bidder do not materially change the allocation, for which the competing bids could be viewed as independent of the strategies submitted by the bidder. In Section 6.1, we extend our analysis for an adaptive adversary, [^12] Also called as non-oblivious adversary (Cesa-Bianchi and Lugosi, 2006) or reactive adversary (Maillard and Munos, 2011). where the competing bids in any round may depend on the bidder’s strategies in the previous rounds.
RoI Constraints in Repeated Setting. In the repeated setting, we require the bidders to satisfy the RoI constraint described in Eq. 2 in every round. A bidding strategy $\mathbf{b}$ is called feasible for a sequence of competing bids $\bm{\beta}_{-}^{t}]_{t\in[T]}$, if the RoI constraint is satisfied for every round $t\in[T]$.
@@
Our notion of RoI constraints in the repeated setting aligns with the definitions in Wilkens et al. (2016, 2017); Lv et al. (2023). Similar constraints (up to scaling factors) are considered by Lucier et al. (2024) and Gaitonde et al. (2023), which they term as maximum bid constraints. [^13] An earlier version of Lucier et al. (2024) termed this as marginal RoI (or value) constraints. Unlike the aggregate constraints over $T$ rounds typically assumed in online ad auction literature (Deng et al., 2021, 2024; Feng et al., 2023; Deng et al., 2023a), we enforce RoI constraints for each auction individually. This distinction reflects the fundamental differences between the two settings: ad auctions often occur simultaneously and frequently, with values in the order of cents, whereas Treasury and emission permit auctions occur sequentially over longer horizons, with units valued in millions. Furthermore, we consider a prior-free setting that allows for strategic (possibly adversarial) behavior by the competitors. Hence, bidders in such auctions prioritize profitability in each auction rather than waiting for an indefinite period, leading to per-round RoI constraints. [^14] Although EU ETS emission permit auctions are scheduled to occur regularly, regulations stipulate that an auction may be canceled if the bidders’ demand falls short of the supply of permits or if the clearing price of the auction does not meet the reserve price (Regulations, 2019). Hence, the bidders are more likely to ensure RoI feasibility in each round. Nonetheless, in Section 6.2, we present a heuristic for the setting where the RoI constraints are enforced only in aggregate over a sliding window of $T_{0}\geq 1$ rounds.
@@ 2.3. Objective and Performance Metric
We consider an online setting where a bidder seeks to maximize cumulative value while satisfying RoI constraints and managing uncertainty about competing bids. In this setting, bidders privately submit their bids, and the auctioneer allocates items and sets prices based on these bids. In round $t$, the learning algorithm maps the feedback information set from the first t1t-1 rounds, denoted by $\mathcal{I}^{t-1}$, to a bid $\mathbf{b}^{t}$, where this mapping may be deterministic or random. Under full information feedback, the information set $\mathcal{I}^{t-1}=(\bm{\beta}^{1},\bm{\beta}^{2},\ldots,\bm{\beta}^{t-1}),$ which includes all bids from previous rounds. In the bandit setting, the information set $\mathcal{I}^{t-1}=(x(\bm{\beta}^{1}),P(\bm{\beta}^{1}),\ldots,x(\bm{\beta}^{t-1}),P(\bm{\beta}^{t-1}))$ includes only the allocations and prices of the previous rounds.
@@
$\textsf{REG}=\max_{\mathbf{b}\in\mathscr{B}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})-\sum_{t=1}^{T}\mathbb{E}[V(\mathbf{b}^{t};\bm{\beta}_{-}^{t})],$ (3)
where the expectation is with respect to any randomness in the learning algorithm. Here, $\mathscr{B}$ is the class of bidding strategies (formally characterized in Section 3). We require the bidding strategies in $\mathscr{B}$ to be RoI feasible for the sequence of competing bids $\bm{\beta}_{-}^{t}]_{t\in[T]}$. In Section 4, we consider the case when both the clairvoyant and the learner choose strategies from the same class and aim to obtain sublinear regret. Later, we consider more challenging settings where the clairvoyant can choose the optimal strategy from much richer classes of strategies compared to the learner. In these cases, we obtain sublinear $\alpha$-approximate regret, where $\alpha\in(0,1]$ is the richness ratio (see Section 5).